# SandCell: Sandboxing Rust Beyond Unsafe Code

### v1

@@ 1 Introduction
Rust is a modern system programming language that provides memory-safety by enforcing strict ownership and borrowing rules at compile time. However, in practice, not all Rust programs fully benefit from compiler-enforced memory safety. For example, pre-compiled Rust libraries or libraries written in other programming languages must be called via foreign-function interfaces (FFIs) which the compiler cannot analyze. Also, because the analysis is conservative, some valid, memory-safe Rust programs will fail to type check. Rust provides an unsafe language superset that allows an “escape hatch" from the compiler-enforced safety analysis.
@@
However, unsafe code is generally hard to get right and has become a common source of bugs in Rust [56]. To write correct unsafe code, programmers need to manually maintain the Rust ownership and borrowing rules and other safety invariants expected by the Rust compiler in the unsafe code block. For example, Rust developers must correctly annotate the lifetime of objects beyond the FFI boundary, and ensure any necessary preconditions necessary for memory-safe execution of C functions are met. Failure to meet these expectations can lead to memory-safety vulnerabilities [5, 30].
@@
Even without unsafe code, safe Rust is not immune to memory-safety issues. For example, a long-existing soundness bug [4] in the type checking of rustc (the official compiler of Rust) allows programmers to arbitrarily expand the lifetime of a reference in safe Rust, which can lead to dangling references. Similarly, the Rust standard library is generally considered part of the safe Rust ecosystem. But given the common use of unsafe code in the standard library, it might contain memory vulnerabilities as well [1, 8, 9].
@@
Given these challenges, there is interest in mitigating lingering memory-safety issues by compartmentalizing Rust programs. A natural idea is to isolate unsafe code from safe code to prevent memory vulnerabilities from affecting the safe subset of the program [18, 53, 39, 48, 20]. These approaches categorize memory objects into safe and unsafe categories according to how they are accessed. Unsafe objects are then placed in a separate compartment, preventing bugs in unsafe code from directly affecting safe objects.
@@
These approaches are, however, limited in several ways. First, their isolation boundary is fixed between safe and unsafe code. In some scenarios discussed in §˜3, this boundary is not expressive enough to accommodate fit-for-purpose isolation policies. Second, these approaches do not support separate sandbox instances for different executions of the same piece of code; all executions of the code would execute within one sandbox and share its privileges and runtime data. As a result, they would not support, for example, isolation between handling different requests in a web server, where a separate sandbox is needed for each request, so one request’s handler cannot interfere with the handlers of other requests.
@@
To address these shortcomings of state-of-the-art compartmentalization solutions for Rust, we propose SandCell, a compiler-based approach with three novel properties. Firstly, to allow flexible isolation boundaries while minimizing developers’ effort on annotation, SandCell leverages syntactic boundaries that already exist in the program (e.g., functions and libraries, called “crates" in Rust  [^1] Rust code is distributed as crates, similar to packages or libraries in other languages.) to let developers specify functions, types, or crates as sandboxes. We refer to sandboxes derived from such syntactic boundaries as syntactic sandbox instances. SandCell automatically partitions aRust application into distinct compartments according to such syntactic instances and inserts sandbox-switching code to enforce the isolation.
@@
  • A compiler-based analysis allowing for users to specify syntactic instances of sandboxes for fine-grained and automated sandboxing of Rust functions, types, modules, and entire crates. We discuss the motivation and challenges in §˜3 and the design in §˜5.
  • Run-time instance sandboxing, which isolates different executions of the same code from each other. We motivate the need in §˜3 and describe the design in §˜5.
@@
  • We evaluate SandCell on several real-world Rust applications, demonstrating its ability to mitigate vulnerabilities and expressing various isolation patterns with reasonable overheads in §˜8,.
2 Background
@@
2.1 Memory Safety in Rust
Safe Rust guarantees memory safety by enforcing strict ownership and borrowing rules at compile time. Informally, it has three rules: 1. Every value has a single owner; assignment to a new owner invalidates the previous owner. 2. At any time, a value can be either modified through one reference (either by the owner or through a single mutable reference) or simultaneously read through multiple immutable references, but not both. 3. No reference can outlive the value it points to. To enforce the third rule, Rust associates a lifetime with each reference at the type-system level, which can be viewed as a set of program points where the reference is valid. The Rust compiler can infer the lifetimes of most references, and the rest must be explicitly annotated by the programmer.
@@ 2.2 Unsafe Rust
In Rust, the unsafe keyword serves two high-level purposes. First, annotating a function definition with unsafe indicates that the function has contracts that the Rust compiler cannot verify. Second, annotating a function call or code block with unsafe indicates that the programmer has manually ensured the contract is satisfied. In practice, programmers use unsafe code for different reasons, and here we discuss the two most common ones [19].
@@
Around 90% of the unsafe code in Rust is used to call unsafe functions [19]. For example, calling a version of array indexing that does not perform bounds checking is unsafe, since the programmer is responsible for ensuring that the index is within bounds. Another example is calling a function from an external library written in C. In these cases, the developer must ensure that preconditions expected by the unsafe function are satisfied. Violating these preconditions can lead to bugs such as RUSTSEC-2023-0044 [30], which is a vulnerability from rust-openssl where a string from Rust code is passed to C code without null-termination, causing an out-of-bound read.
@@
Another common use case of unsafe code is to manipulate raw pointers. It might cause bugs when, for example, the developer assigns an incorrect lifetime to a raw pointer that is cast to a reference. Since the lifetime of a raw pointer is not tracked by the compiler, such bugs cannot be detected at compile time. As a practical example, CVE-2025-24898 [5] is a vulnerability in rust-openssl where an incorrect lifetime parameter is bound to a buffer (allocated in the C code). Because the lifetime parameter exceeds the actual lifetime of the buffer, it leads to a use-after-free vulnerability.
2.3 Bugs in Safe Rust
@@
The official Rust compiler, rustc, is not bug-free either and might accept programs that violate its safety invariants. For example, a well-known bug [17] in the type checker enables the lifetime of a local reference to be expanded to the global lifetime. Exploiting such soundness bugs, adversaries may intentionally upload vulnerable crates to crates.io, the Rust package index, to mount supply-chain attacks. One tool Rust developers have at their disposal to minimize the risk posed by third-party dependencies is limiting the dependencies to crates without unsafe code. The package index tags such crates with the “no_unsafe" marker to make finding such crates easier. This mechanism incentivizes malicious developers to exploit soundness bugs instead of unsafe code. The rustc team maintains a list of such compiler bugs [10].
@@
The Rust standard library is also an important building block for safe Rust code as it provides safe abstractions for low-level operations such as heap allocation and synchronization primitives. But bugs in the standard library also exist. For example, an integer overflow led to a buffer overflow in std::str::repeat [12].
2.4 In-process isolation
@@
One way to sandbox cooperating software modules is to place each in a process with its own virtual address space [42]. However, for tightly-coupled modules, process-based isolation incurs prohibitive context switch and communication overhead [34]. In-process isolation is a security technique that places the code and data for a sandboxed module in its own fault domain, a logically separate portion of the application’s address space. In-process isolation trades strong, hardware- and OS-enforced process isolation for substantially faster context switches between fault domains. In-process isolation can be realized portably and language-independently through software-based fault isolation (SFI) [62, 59]. SFI, however, increases execution times for the sandboxed modules as their memory accesses must be checked for validity by a software-enforced in-line reference monitor. Modern approaches for in-process isolation leverage hardware features, such as memory protection keys (MPK), to improve upon the efficiency of SFI. The popularity of sandboxed JavaScript in modern browsers, managed languages, and application sandboxing in mobile OSes have motivated processor manufacturers to include such hardware features in commercial, off-the-shelf processors.
@@
In contrast to OS-enforced paging, the PKRU configuration can be modified by the userspace process. Since PKU hardware enforces access control based on the current configuration of the PKRU, unauthorized modification of the PKRU, or techniques that bypass PKU enforcement completely makes PKU vulnerable to threats from run-time attacks and malicious software aiming to escape from its sandbox. For example, an attacker could manipulate /proc/self/mem to bypass MPK protections [24] or misuse the wrpkru instruction to modify the PKRU register [60].
@@
Figure 1: A simplified web server. In Rust, modules (the keyword mod) are used to organize definitions such as functions and types.
3 Motivation and Challenges
Given the disposition of vulnerabilities in unsafe Rust code to undermine the compiler-enforced memory-safety properties, a natural idea is to isolate unsafe code from the rest of a Rust program, as explored by previous work [18, 53, 39, 48, 20]. However, simply having two compartments (one safe and one unsafe compartment) is insufficient for the reasons discussed in §˜1 and §˜2. Consequently, compartmentalization in Rust must solve the following challenges.
@@ 3 Motivation and Challenges
The Rust compiler itself is not bug free and might accept programs that violate its safety invariants, as discussed in §˜2.3. By sandboxing untrusted Rust code, the impact of such bugs can be limited to the sandboxed code and data. Such cases requires the ability to isolate a mix of safe and unsafe code.
Rust’s standard library is another important building block for its safety guarantees and provides safe abstractions for low-level operations such as heap allocation and synchronization primitives. Previous approaches that isolate unsafe code from safe code generally include the standard library in the TCB. To mitigate the impact of bugs in the Rust standard library (e.g., [12]), the compartmentalization method must support sandboxing not only untrusted and user-specified code, but also dependencies and their transitive dependencies.
@@
4 System and Adversary Model
@@
We assume an adversary who exploits either pre-existing vulnerabilities in Rust applications, or intentionally introduces vulnerabilities they can exploit into crates that they control. Using such vulnerabilities, the adversary can mount control-flow or data-only attacks against the victim application. The adversary may also attempt to misuse privileges granted to the application, such as performing unwanted, harmful system calls. The goal of SandCell is to mitigate such attacks by isolating untrusted code and data from the rest of the program. However, we do not provide security guarantees against attackers specifically targeting the user-space system-call interpositioning. It is also possible that the attacker can indirectly impact the program through information flow between trusted and untrusted code such as through side channels, which is not in the scope of SandCell.
5 Design
In this section, we give an overview of the design of SandCell, including the overall workflow, user specification, its semantics, and the run-time architecture. We will delay the discussion of the program analysis supporting SandCell until §˜6 and the implementation details until §˜7.
5.1 Overview
SandCell takes the Rust source code and a sandboxing specification as input, and produces a sandboxed binary as output. The SandCell specification contains a list of functions, types, or crates to be sandboxed. We will elaborate on this in §˜5.2. SandCell’s analysis builds a call graph of the program and uses the specification to identify the instrumentation points necessary for sandboxing the specified components. SandCell also uses flow-insensitive data-flow analysis to identify allocation sites that might allocate heap objects that are accessed across sandbox boundaries (§˜6). These results are used by the instrumentation component (§˜7.2) of SandCell to alter the program’s source code and insert calls into the low-level isolation application programming interfaces (API) at the instrumentation points identified during the analysis. The in-process isolation APIs and the run-time architecture of SandCell will be described in §˜5.4. Finally, the instrumented program is compiled by rustc as normal, producing a sandboxed binary.
5.2 User Specification
@@
Figure 2: A SandCell specification example.
In this section, we describe the syntax and semantics of the SandCell specification and how they address the challenges discussed in §˜3.
To address Challenge I, SandCell allows users to define sandboxes for functions, modules, crates, and types by listing their names in the specification file, as illustrated by the example in Figure˜2, which specifies that functions foo and bar should be sandboxed. When there are multiple candidates with the same name, users can refine it with quantified paths.
@@
Invocations of different units (functions, crates, etc.) create different sandboxes. For the example in Figure 2, invocations of functions foo and bar create different sandboxes; so their executions are isolated from each other. Conceptually, a sandbox is a piece of memory region that contains the code and data of the sandboxed unit (function, crate, etc.) as well as its transitive dependencies. For example, if both foo and bar calls another function baz, there is a separate copy of the code and data of baz in each sandbox.
@@
In SandCell, programmers specify sandbox boundaries using syntactic boundaries of program units such as functions and crates. There are other alternative ways of specifying sandbox boundaries. For example, tools such as Privtrans [21], SeCage [50], and PtrSplit [49] allow users to specify sensitive data and automatically computes a sandbox boundary using static analysis. We find that in our use cases the trust boundary is often aligned with the syntactic boundaries. Also, the syntactic based method is more convenient to use without knowing too much detail of the program about what data is considered sensitive.
5.3 Analysis and Instrumentation

### v2

@@ 1. Introduction
Rust is a modern system programming language that provides memory-safety by enforcing strict ownership and borrowing rules at compile time. However, in practice, not all Rust programs fully benefit from compiler-enforced memory safety. For example, pre-compiled Rust libraries or libraries written in other programming languages must be called via foreign-function interfaces (FFIs) which the compiler cannot analyze. Also, because the analysis is conservative, some valid, memory-safe Rust programs will fail to type check. Rust provides an unsafe language superset that allows an “escape hatch” from the compiler-enforced safety analysis.
@@
However, unsafe code is generally hard to get right and has become a common source of bugs in Rust (rustfuzztrophy, ). To write correct unsafe code, programmers need to manually maintain the Rust ownership and borrowing rules and other safety invariants expected by the Rust compiler in the unsafe code block. For example, Rust developers must correctly annotate the lifetime of objects beyond the FFI boundary, and ensure any necessary preconditions necessary for memory-safe execution of C functions are met. Failure to meet these expectations can lead to memory-safety vulnerabilities (nistCVE202524898, ; rustsecRUSTSEC20230044Openssl, ).
@@
Even without unsafe code, safe Rust is not immune to memory-safety issues. For example, a long-existing soundness bug (rust-implied-bounds-bug, ) in the type checking of rustc (the official compiler of Rust) allows programmers to arbitrarily expand the lifetime of a reference in safe Rust, which can lead to dangling references. Similarly, the Rust standard library is generally considered part of the safe Rust ecosystem. But given the common use of unsafe code in the standard library, it might contain memory vulnerabilities as well (githubBTreeSetCauses, ; githubFaultPushing, ; githubSoundnessIssue, ).
@@
Given these challenges, there is interest in mitigating lingering memory-safety issues by compartmentalizing Rust programs. A natural idea is to isolate unsafe code from safe code to prevent memory vulnerabilities from affecting the safe subset of the program (almohriFideliusCharmIsolating2018, ; riveraKeepingSafeRust2021, ; kirthPKRUsafeAutomaticallyLocking2022, ; liuSecuringUnsafeRust2020, ; bangTRustCompilationFramework2023, ). These approaches categorize memory objects into safe and unsafe categories according to how they are accessed. Unsafe objects are then placed in a separate compartment, preventing bugs in unsafe code from directly affecting safe objects.
@@
These approaches are, however, limited in several ways. First, their isolation boundary is fixed between safe and unsafe code. In some scenarios discussed in § 3, this boundary is not expressive enough to accommodate fit-for-purpose isolation policies. Second, these approaches do not support separate sandbox instances for different executions of the same piece of code; all executions of the code would execute within one sandbox and share its privileges and runtime data. As a result, they would not support, for example, isolation between handling different requests in a web server, where a separate sandbox is needed for each request, so one request’s handler cannot interfere with the handlers of other requests.
@@
To address these shortcomings of state-of-the-art compartmentalization solutions for Rust, we propose SandCell, a compiler-based approach with three novel properties. Firstly, to allow flexible isolation boundaries while minimizing developers’ effort on annotation, SandCell leverages syntactic boundaries that already exist in the program (e.g., functions and libraries, called “crates” in Rust  [^1] Rust code is distributed as crates, similar to packages or libraries in other languages.) to let developers specify functions, types, or crates as sandboxes. We refer to sandboxes derived from such syntactic boundaries as syntactic sandbox instances. SandCell automatically partitions a Rust application into distinct compartments according to such syntactic instances and inserts sandbox-switching code to enforce the isolation.
@@
  • A compiler-based analysis allowing for users to specify syntactic instances of sandboxes for fine-grained and automated sandboxing of Rust functions, types, modules, and entire crates. We discuss the motivation and challenges in § 3 and the design in § 5.
  • Run-time instance sandboxing, which isolates different executions of the same code from each other. We motivate the need in § 3 and describe the design in § 5.
@@
  • We evaluate SandCell on several real-world Rust applications, demonstrating its ability to mitigate vulnerabilities and expressing various isolation patterns with reasonable overheads in § 8.
2. Background
@@
2.1. Memory Safety in Rust
Safe Rust guarantees memory safety by enforcing strict ownership and borrowing rules at compile time. Informally, it has three rules: (1) Every value has a single owner; assignment to a new owner invalidates the previous owner. (2) At any time, a value can be either modified through one reference (either by the owner or through a single mutable reference) or simultaneously read through multiple immutable references, but not both. (3) No reference can outlive the value it points to. To enforce the third rule, Rust associates a lifetime with each reference at the type-system level, which can be viewed as a set of program points where the reference is valid. The Rust compiler can infer the lifetimes of most references, and the rest must be explicitly annotated by the programmer.
@@ 2.2. Unsafe Rust
In Rust, the unsafe keyword serves two high-level purposes. First, annotating a function definition with unsafe indicates that the function has contracts that the Rust compiler cannot verify. Second, annotating a function call or code block with unsafe indicates that the programmer has manually ensured the contract is satisfied. In practice, programmers use unsafe code for different reasons, and here we discuss the two most common ones (astrauskasHowProgrammersUse2020, ).
@@
Around 90% of the unsafe code in Rust is used to call unsafe functions (astrauskasHowProgrammersUse2020, ). For example, calling a version of array indexing that does not perform bounds checking is unsafe, since the programmer is responsible for ensuring that the index is within bounds. Another example is calling a function from an external library written in C. In these cases, the developer must ensure that preconditions expected by the unsafe function are satisfied. Violating these preconditions can lead to bugs such as RUSTSEC-2023-0044 (rustsecRUSTSEC20230044Openssl, ), which is a vulnerability from rust-openssl where a string from Rust code is passed to C code without null-termination, causing an out-of-bound read.
@@
Another common use case of unsafe code is to manipulate raw pointers. It might cause bugs when, for example, the developer assigns an incorrect lifetime to a raw pointer that is cast to a reference. Since the lifetime of a raw pointer is not tracked by the compiler, such bugs cannot be detected at compile time. As a practical example, CVE-2025-24898 (nistCVE202524898, ) is a vulnerability in rust-openssl where an incorrect lifetime parameter is bound to a buffer (allocated in the C code). Because the lifetime parameter exceeds the actual lifetime of the buffer, it leads to a use-after-free vulnerability.
2.3. Bugs in Safe Rust
@@
The official Rust compiler, rustc, is not bug-free either and might accept programs that violate its safety invariants. For example, a well-known bug (rust25860, ) in the type checker enables the lifetime of a local reference to be expanded to the global lifetime. Exploiting such soundness bugs, adversaries may intentionally upload vulnerable crates to crates.io, the Rust package index, to mount supply-chain attacks. One tool Rust developers have at their disposal to minimize the risk posed by third-party dependencies is limiting the dependencies to crates without unsafe code. The package index tags such crates with the “no_unsafe” marker to make finding such crates easier. This mechanism incentivizes malicious developers to exploit soundness bugs instead of unsafe code. The rustc team maintains a list of such compiler bugs (rust-unsound-bug-list, ).
@@
The Rust standard library is also an important building block for safe Rust code as it provides safe abstractions for low-level operations such as heap allocation and synchronization primitives. But bugs in the standard library also exist. For example, an integer overflow led to a buffer overflow in std::str::repeat (RustStdCVE20181000810, ).
2.4. In-process isolation
@@
One way to sandbox cooperating software modules is to place each in a process with its own virtual address space (lamowskiSandcrustAutomaticSandboxing2017, ). However, for tightly-coupled modules, process-based isolation incurs prohibitive context switch and communication overhead (gulmezFriendFoeExploring2023, ). In-process isolation is a security technique that places the code and data for a sandboxed module in its own fault domain, a logically separate portion of the application’s address space. In-process isolation trades strong, hardware- and OS-enforced process isolation for substantially faster context switches between fault domains. In-process isolation can be realized portably and language-independently through software-based fault isolation (SFI) (wahbe1993efficient, ; Tan17, ). SFI, however, increases execution times for the sandboxed modules as their memory accesses must be checked for validity by a software-enforced in-line reference monitor. Modern approaches for in-process isolation leverage hardware features, such as memory protection keys (MPK), to improve upon the efficiency of SFI. The popularity of sandboxed JavaScript in modern browsers, managed languages, and application sandboxing in mobile OSes have motivated processor manufacturers to include such hardware features in commercial, off-the-shelf processors.
@@
In contrast to OS-enforced paging, the PKRU configuration can be modified by the userspace process. Since PKU hardware enforces access control based on the current configuration of the PKRU, unauthorized modification of the PKRU, or techniques that bypass PKU enforcement completely makes PKU vulnerable to threats from run-time attacks and malicious software aiming to escape from its sandbox. For example, an attacker could manipulate /proc/self/mem to bypass MPK protections (connor20, ) or misuse the wrpkru instruction to modify the PKRU register (Vahldiek-Oberwagner19, ).
@@
Figure 1. A simplified web server. In Rust, modules (the keyword mod) are used to organize definitions such as functions and types.
3. Motivation and Challenges
Given the disposition of vulnerabilities in unsafe Rust code to undermine the compiler-enforced memory-safety properties, a natural idea is to isolate unsafe code from the rest of a Rust program, as explored by previous work (almohriFideliusCharmIsolating2018, ; riveraKeepingSafeRust2021, ; kirthPKRUsafeAutomaticallyLocking2022, ; liuSecuringUnsafeRust2020, ; bangTRustCompilationFramework2023, ). However, simply having two compartments (one safe and one unsafe compartment) is insufficient for the reasons discussed in § 1 and § 2. Consequently, compartmentalization in Rust must solve the following challenges.
@@ 3. Motivation and Challenges
The Rust compiler itself is not bug free and might accept programs that violate its safety invariants, as discussed in § 2.3. By sandboxing untrusted Rust code, the impact of such bugs can be limited to the sandboxed code and data. Such cases requires the ability to isolate a mix of safe and unsafe code.
Rust’s standard library is another important building block for its safety guarantees and provides safe abstractions for low-level operations such as heap allocation and synchronization primitives. Previous approaches that isolate unsafe code from safe code generally include the standard library in the TCB. To mitigate the impact of bugs in the Rust standard library (e.g., (RustStdCVE20181000810, )), the compartmentalization method must support sandboxing not only untrusted and user-specified code, but also dependencies and their transitive dependencies.
@@
4. System and Adversary Model
@@
We assume an adversary who exploits either pre-existing vulnerabilities in Rust applications, or intentionally introduces vulnerabilities they can exploit into crates that they control. Using such vulnerabilities, the adversary can mount control-flow or data-only attacks against the victim application. The adversary may also attempt to misuse privileges granted to the application, such as performing unwanted, harmful system calls. The goal of SandCell is to mitigate such attacks by isolating untrusted code and data from the rest of the program. However, we do not provide security guarantees against attackers specifically targeting the user-space system-call interpositioning (see  § 7.4). It is also possible that the attacker can indirectly impact the program through information flow between trusted and untrusted code such as through side channels, which is not in the scope of SandCell.
5. Design
In this section, we give an overview of the design of SandCell, including the overall workflow, user specification, its semantics, and the run-time architecture. We will delay the discussion of the program analysis supporting SandCell until § 6 and the implementation details until § 7.
5.1. Overview
SandCell takes the Rust source code and a sandboxing specification as input, and produces a sandboxed binary as output. The SandCell specification contains a list of functions, types, or crates to be sandboxed. We will elaborate on this in § 5.2. SandCell’s analysis builds a call graph of the program and uses the specification to identify the instrumentation points necessary for sandboxing the specified components. SandCell also uses flow-insensitive data-flow analysis to identify allocation sites that might allocate heap objects that are accessed across sandbox boundaries (§ 6). These results are used by the instrumentation component (§ 7.2) of SandCell to alter the program’s source code and insert calls into the low-level isolation application programming interfaces (API) at the instrumentation points identified during the analysis. The in-process isolation APIs and the run-time architecture of SandCell will be described in § 5.4. Finally, the instrumented program is compiled by rustc as normal, producing a sandboxed binary.
5.2. User Specification
@@
Figure 2. A SandCell specification example.
In this section, we describe the syntax and semantics of the SandCell specification and how they address the challenges discussed in § 3.
To address Challenge I, SandCell allows users to define sandboxes for functions, modules, crates, and types by listing their names in the specification file, as illustrated by the example in Figure 2, which specifies that functions foo and bar should be sandboxed. When there are multiple candidates with the same name, users can refine it with quantified paths.
@@
Invocations of different units (functions, crates, etc.) create different sandboxes. For the example in Figure 2, invocations of functions foo and bar create different sandboxes; so their executions are isolated from each other. Conceptually, a sandbox is a memory region that contains the code and data of the sandboxed unit (function, crate, etc.) as well as its transitive dependencies. For example, if both foo and bar calls another function baz, there is a separate copy of the code and data of baz in each sandbox.
@@
Specification strategies.
In SandCell, programmers specify sandbox boundaries using syntactic boundaries of program units such as functions and crates. Identifying good isolation boundaries is a well-studied problem: tools such as Privtrans (brumley2004privtrans, ), SeCage (liu2015thwarting, ), PtrSplit (liu2017ptrsplit, ), and Program-mandering (programMandering, ) can automatically compute sandbox boundaries based on information flow analysis or sensitive data annotations. However, these automatic approaches typically enforce fixed kinds of policies (e.g., protecting confidential information). On the other hand, developers with deep knowledge of a codebase understand the program’s behavior and trust relationship and can thus develop policies targeting custom security properties.
Since SandCell is orthogonal to both approaches, it can be combined with automatic partitioning tools or expert-crafted policies. However, even without such tools or expertise, developers can follow simple strategies that provide meaningful security benefits:
  • Library-based (supply-chain): Sandbox entire third-party libraries, especially those with known vulnerabilities or written in unsafe languages. This strategy mitigates supply-chain attacks without requiring understanding of the library’s internals.
  • User-based: In multi-tenant applications, isolate components that process different users’ data from each other. This prevents a compromised user session from attacking other users.
  • Input-oriented: Sandbox components that directly handle external input, such as parsers, decoders, and deserializers. These components are common attack surfaces and benefit from isolation regardless of whether specific vulnerabilities are known.
In our case studies, we found that trust boundaries align naturally with syntactic boundaries (crates, modules, functions), making specification straightforward. We evaluate these strategies in § 8.
5.3. Analysis and Instrumentation